OpenStudy (anonymous):

lim as x approaches pi/4: [(tan x - 1)/(x - pi/4)] The limit represents the derivative of some function f at some number a. State such an f and a.

OpenStudy (anonymous):

this is the derivative of guess what?

OpenStudy (anonymous):

it is the derivative of tangent, at \(\frac{\pi}{4}\)

OpenStudy (anonymous):

\[\frac{f(x)-f(a)}{x-a}\] if \(f(x)=\tan(x), a=\frac{\pi}{4}\)